Update 24 September 2016
Full magnet string
Please join me in congratulating all involved in bringing the MICE magnets to current in a configuration typical of the settings required to deliver the field-on scattering programme at Step IV. The magnets have been at current since the evening of Thursday 22Sep16 and a full 48-hour soak test is now in progress. On Friday 23Sep16, the the first data was taken in this configuration with 140 MeV and 170 beams, i.e. we’ve started the empty-channel data taking for the field-on scattering programme.